OpenStudy (anonymous):

a student can type 180 words in 3 min. and 368 words in 7 min. how many words can the student type in 10 minutes?

OpenStudy (dumbcow):

slope = (368-180)/(7-3) slope = 188/4 = 47 wpm words= 368 + 47(3) =368+141 =509